use Chebyshevs theorem for disturbution with a mean of 80 and sd at 10. the question is what percentage of values will fall bwtn 65 and 95 %.
this is my answer.
80 +3(10)=80 + 30 =110 80 - 3(10)= 80 - 50 = 30 I said the percentage was 88.89%. I am lost in regards to this problem. Could you show me step by step what to do.

Answer by MathTherapy(10552)

Probability-and-statistics/176325 (2009-01-12 23:36:55): use Chebyshevs theorem for disturbution with a mean of 80 and sd at 10. the question is what percentage of values will fall bwtn 65 and 95 %.
this is my answer.
80 +3(10)=80 + 30 =110 80 - 3(10)= 80 - 50 = 30 I said the percentage was 88.89%. I am lost in regards to this problem. Could you show me step by step what to do.

You're really lost. lol.

Data between 65 and 95% (95.44% to be exact) represent a total of 2 standard deviations from the mean. According to Chebyshev's theorem, at least 1 - 1/k^2 of data values will lie within k standard deviations of the mean (DeSanto & Totoro, 1995, p. 171), and where k > 1.
Therefore, 1 - 1/k^2 = 1 - 1/2^2 = 1 - 1/4 = 3/4, or 75%.